# Comparison of Extended Lubrication Theories for Stokes Flow

### v3

@@ 1 Introduction
The classical lubrication assumptions of a long and thin fluid domain and a small scaled Reynolds number are used to justify neglecting the inertial velocity and cross-film pressure gradient terms in the Navier-Stokes equations. The resulting Reynolds equation is a dramatic simplification of the original governing equations. More recently, variations on the thin film assumption have given rise to higher-order approximations for lubrication theory [venner_stokes_2003, marusic-paloka_second_2012, tavakol_extended_2017, takeuchi_extended_2019, sialmas_newtonian_2024, banerjee_variational_2024]. These extended lubrication theory models retain certain simplifying assumptions as in classical lubrication theory, while nevertheless being applicable to a broader scope of geometries in the low Reynolds number limit.
@@
Previous research on the flows modeled by lubrication theory has highlighted the sensitivity of lubrication theory to variations of the fluid film height. In particular, in the limit of zero Reynolds number flows, discrepancies between lubrication theory and Stokes flow increase with both the magnitude and frequency of large gradients in the height [brown_applicability_1995, biswas_backward-facing_2004, dobrica_about_2009]. The pressure drop modeled by the Reynolds equation may be underestimated due to additional pressure losses at points of sudden expansion that are not accurately modeled [dobrica_reynolds_2005, dobrica_about_2009]. Furthermore, notable flow features of corner recirculation seen in the Stokes solution are not accurately captured with the Reynolds equation [armaly_experimental_1983, biswas_backward-facing_2004, shyu_numerical_2018].
@@
In this paper, we consider classical lubrication theory (the Reynolds equation), and more recent variations on lubrication theory which we refer to as extended lubrication theory [takeuchi_extended_2019] and/or perturbed lubrication theory [venner_stokes_2003, marusic-paloka_second_2012, tavakol_extended_2017, sialmas_newtonian_2024] in the case of models based on perturbation theory. We asses the performance of these models at low Reynolds numbers by comparison to the Stokes equations. We propose a formulation of extended lubrication theory that builds on the model of Takeuchi & Gu (2019) [takeuchi_extended_2019]; our approach modifies the treatment of the velocity boundary conditions and ensures incompressibility. The pressure and velocity solutions from each model are compared in a variety of geometries: a logistic step, a backward facing step, and two versions of a triangular slider that we refer to as positive and negative based on the texture to be defined later on. Through these examples, we assess the sensitivity of each model to varying magnitudes of surface variation and discontinuities in the surface gradient in the low Reynolds number limit.
@@ 2.1 Lubrication Theory
We begin with a review of classical lubrication theory. This treatment may be found in various textbooks [leal_advanced_2007], and we include it here to establish notation and to make the presentation self-contained. The Navier-Stokes equations,
(1) $\frac{\partial p}{\partial x}$ $=\eta\Big(\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{2}u}{\partial y^{2}}\Big)-\rho\Big(u\frac{\partial u}{\partial x}+v\frac{\partial u}{\partial y}\Big),$
(2) $\frac{\partial p}{\partial y}$ $=\eta\Big(\frac{\partial^{2}v}{\partial x^{2}}+\frac{\partial^{2}v}{\partial y^{2}}\Big)-\rho\Big(u\frac{\partial v}{\partial x}+v\frac{\partial v}{\partial y}\Big),$
@@
Lubrication theory provides a linearized approximation to the Navier-Stokes equations valid for geometries in which the fluid is constrained to a relatively thin domain. For the two-dimensional fluid domain $(x,y)\in[x_{0},x_{f}]\times[0,h(x)]$, the characteristic length scales are $L_{x}=x_{f}-x_{0}$ and $L_{y}=\max h(x)>0$. Given a prescribed constant flux $\mathcal{Q}\neq 0$, the characteristic velocities are $U_{*}=\mathcal{Q}/L_{y}$ and $V_{*}=\mathcal{Q}/L_{x}$, and the characteristic pressure is $P_{*}=\eta U_{*}L_{x}/L_{y}^{2}$. The dimensionless variables are $X=x/L_{x}$, $Y=y/L_{y}$, $H=h/L_{y}$, $U=u/U_{*}$, $V=v/V_{*}$, and $P=p/P_{*}$. The Reynolds number is $\text{Re}=\rho U_{*}L_{x}/\eta$ and the length scale ratio is $\varepsilon=L_{y}/L_{x}$. The Navier-Stokes equations Eqs. 1 and 2 in dimensionless terms are,
(4) $\frac{\partial P}{\partial X}$ $=\Big(\varepsilon^{2}\frac{\partial^{2}U}{\partial X^{2}}+\frac{\partial^{2}U}{\partial Y^{2}}\Big)-\varepsilon^{2}\text{Re}\Big(U\frac{\partial U}{\partial X}+V\frac{\partial U}{\partial Y}\Big),$
(5) $\frac{\partial P}{\partial Y}$ $=\varepsilon^{2}\Big(\varepsilon^{2}\frac{\partial^{2}V}{\partial X^{2}}+\frac{\partial^{2}V}{\partial Y^{2}}\Big)-\varepsilon^{4}\text{Re}\Big(U\frac{\partial V}{\partial X}+V\frac{\partial V}{\partial Y}\Big),$
@@
We also prescribe the fixed outlet pressure $P(1,0)=0$. Because the flux is prescribed, the average pressure drop $\Delta P$ is given from the solution, as in [banerjee_variational_2024],
(9) $\Delta P=\frac{1}{H(0)}\int_{0}^{H(0)}P(0,Y)\,dY-\frac{1}{H(1)}\int_{0}^{H(1)}P(1,Y)\,dY.$
Conversely, one may prescribe the average pressure drop, in which case the flux would be determined by the solution.
@@
where $P_{0}$, $U_{0}$, $V_{0}$ denote the order $\varepsilon^{0}$ lubrication approximation. Together with incompressibility Eq. 3, Eqs. 10 and 11 constitute the governing equations for lubrication theory.
@@
Note that the flux appears in the scaling of the boundary velocity, $U_{b}=\mathcal{U}L_{y}/\mathcal{Q}$. Since the flux is prescribed, the inlet pressure $P_{0}(0)$ and the average pressure drop $\Delta P_{0}$ are determined by the solution. The solution $P_{0}$ to the Reynolds equation is determined using a second-order accurate finite difference scheme, see Appendix A.
@@ 2.2 Perturbed Lubrication Theory
On substituting the ansatz Eqs. 16, 17 and 18 into Eqs. 4 and 5 at order $\varepsilon^{2}$, the velocity and pressure satisfy the following reduced equations,
@@
The order $\varepsilon^{0}$ Reynolds solution $U_{0}$, $V_{0}$, and $P_{0}$ satisfies the boundary conditions Eqs. 6, 7 and 15 respectively. The higher-order velocities $U$ and $V$ also satisfy Eqs. 6 and 7. The higher-order pressures $P$ satisfy $P(1,0)=0$, but do not necessarily satisfy the pressure gradient condition in Eq. 15. Instead, we require that the flux $\mathcal{Q}$ be the same as the flux of the order $\varepsilon^{0}$ solution, such that, in dimensionless terms, $\int_{0}^{H}U_{0}\,dY=1$ and $\int_{0}^{H}U_{k>0}\,dY=0$. Because the flux is prescribed, the average pressure drop $\Delta P_{\varepsilon^{k}\text{-PLT}}$ is determined by the solution. In general, the average pressure drops $\Delta P_{\varepsilon^{2}\text{-PLT}}=\Delta P_{0}+\varepsilon^{2}\Delta P_{2}$ and $\Delta P_{\varepsilon^{4}\text{-PLT}}=\Delta P_{0}+\varepsilon^{2}\Delta P_{2}+\varepsilon^{4}\Delta P_{4}$ will deviate from $\Delta P_{0}$. We determine $\Delta P_{k>0}$ through numerical integration according to Eq. 9.
@@
The method of perturbed solutions is also well suited to techniques such as diagonal Padé approximants or the Shanks’ transform [sialmas_newtonian_2024]. These techniques may be used to accelerate and extend the domain of the convergence for the series solution. Although such numerical methods are out of the scope of the present work, we consider this a potential avenue for future investigation and improvement of the PLT method.
@@ 3 Velocity-Adjusted Extended Lubrication Theory
Observe that if we assume $\varsigma=0$, the velocity field from Eqs. 27 and 28 is not necessarily incompressible, and Eq. 28 may not satisfy $V(X,H)=0$.
@@ 4.1 The logistic step
First, we consider the logistic step, shown in Fig. 1. The height function is given by,
@@
The errors in velocity and pressure compared with the Stokes solution are shown in Fig. 3. For the wide range of $\lambda<32$, the VA-ELT solution of pressure is significantly more accurate than the Reynolds and T.G.-ELT solutions, and the T.G.-ELT solution of pressure yields a marginal improvement over the Reynolds pressure. For $2<\lambda<8$, the $\varepsilon^{2}$-PLT solution is best at approximating the pressure, and the VA-ELT solution is best at approximating the velocity. The $\varepsilon^{2}$-PLT solution performs better than the $\varepsilon^{4}$-PLT solution for all $\lambda>2$. For $\lambda\geq 32$, the Reynolds solution has the smallest error in velocity and pressure; the Reynolds solution has the most stable error across the range of $\lambda$. Overall, both the $\varepsilon^{2}$-PLT and VA-ELT solutions perform well for the logistic step up to moderate surface variation. We note however, that all the lubrication models discussed here break down in the limit $\lambda\to\infty$ of a discontinuous step, see Appendix D.
@@ 4.2 The triangular slider
Next, we consider a triangular slider, shown in Fig. 4. The height function is given by,
@@
The errors in pressure and velocity compared with the Stokes solution for the triangular slider are shown in Fig. 7. The error in velocity is very similar for all of the models. For pressure, the VA-ELT solution is a consistent improvement compared to the Reynolds and T.G.-ELT solutions, and the $\varepsilon^{k}$-PLT solutions perform better still. For negative texturing ($H_{v}<1$), the $\varepsilon^{k}$-PLT solutions of pressure have significantly smaller error relative to both the VA-ELT and Reynolds solutions; for positive texturing ($H_{v}>1$), the $\varepsilon^{k}$-PLT and VA-ELT solutions are more similar. The error in both pressure and velocity decreases as the difference between the minimum and maximum height decreases. In particular, for the case of negative texturing, the increasing magnitude of surface gradient as the minimum height decreases corresponds to an increase in error for pressure and velocity; this emphasizes the importance of considering the magnitude of surface gradients as well as the length scale ratio when selecting a fluid model.
@@ Appendix A The Reynolds equation finite difference solution
Define the uniform discretisation of $0,1]=([x_{0},x_{L}]-x_{0})/L_{x}$,
@@
Appendix B The Stokes Equations
The biharmonic formulation of the incompressible Navier-Stokes equations is an effective method of solution for low Reynolds number flows [gupta_new_2005, marner_potential-velocity_2014, sen_4oec_2015, biswas_hoc_2017]. Compared with the lubrication approximation, the bi-harmonic Navier-Stokes equations are formulated with a single characteristic velocity $U_{*}$ and a single length scale $L_{x}$. The characteristic pressure is then $P_{*}=\eta U_{*}/L_{x}$. Consequently, the biharmonic Navier-Stokes model is not sensitive to surface variation as the lubrication models are.
With the dimensionless variables $X=x/L_{x}$, $Y=y/L_{x}$, $U=u/U_{*}$, $V=v/U_{*}$ and $P=p/P_{*}$, the Navier-Stokes equations Eqs. 1 and 2 are expressed as
(34) $\frac{\partial P}{\partial X}$ $=\Big(\frac{\partial^{2}U}{\partial X^{2}}+\frac{\partial^{2}U}{\partial Y^{2}}\Big)-\text{Re}\Big(U\frac{\partial U}{\partial X}+V\frac{\partial U}{\partial Y}\Big),$
(35) $\frac{\partial P}{\partial Y}$ $=\Big(\frac{\partial^{2}V}{\partial X^{2}}+\frac{\partial^{2}V}{\partial Y^{2}}\Big)-\text{Re}\Big(U\frac{\partial V}{\partial X}+V\frac{\partial V}{\partial Y}\Big).$

### v4

@@ 1 Introduction
The classical lubrication assumptions of a long and thin fluid domain and a small scaled Reynolds number are used to justify neglecting the inertial velocity and cross-film pressure gradient terms in the Navier-Stokes equations. The resulting Reynolds equation is a dramatic simplification of the original governing equations. More recently, variations on the thin film assumption have given rise to higher-order approximations for lubrication theory [venner_stokes_2003, marusic-paloka_second_2012, tavakol_extended_2017, takeuchi_extended_2019, sialmas_newtonian_2024, okazaki_fluid_2024]. These extended lubrication theory models retain certain simplifying assumptions as in classical lubrication theory, while nevertheless being applicable to a broader scope of geometries in the low Reynolds number limit.
@@
Previous research on the flows modeled by lubrication theory has highlighted the sensitivity of lubrication theory to variations of the fluid film height. In particular, in the limit of zero Reynolds number flows, discrepancies between lubrication theory and Stokes flow increase with both the magnitude and frequency of large gradients in the height [brown_applicability_1995, biswas_backward-facing_2004, dobrica_about_2009]. The pressure drop modeled by the Reynolds equation may be underestimated due to additional pressure losses at points of sudden expansion that are not accurately modeled [dobrica_reynolds_2005, dobrica_about_2009]. Furthermore, notable flow features of corner recirculation seen in the Stokes solution are not accurately captured with the Reynolds equation [armaly_experimental_1983, biswas_backward-facing_2004, shyu_numerical_2018]. Note that the references cited above should not be interpreted as an exhaustive review; perturbation approaches are a rich area of research in the field of fluid dynamics. The references [venner_stokes_2003, marusic-paloka_second_2012, tavakol_extended_2017, takeuchi_extended_2019, sialmas_newtonian_2024, okazaki_fluid_2024] focus on Newtonian fluids and perturbation of the length scale ratio which quantifies the thin film assumption of classical lubrication theory. For further discussion of perturbation theories, including for non-Newtonian fluids, we direct interested readers to [tshehla_flow_2013, housiadas_high-order_2016, synder_comparison_2018, ahmed_new_2021, boyko_pressure-driven_2022, abbaspur_analytical_2023, housiadas_pressure-drop_2024, okazaki_fluid_2024, ahmed_viscoelastic_2025, sialmas_general_2025] and references therein.
@@
In this paper, we consider classical lubrication theory (the Reynolds equation), and more recent variations on lubrication theory which we refer to as extended lubrication theory [takeuchi_extended_2019, okazaki_fluid_2024] or perturbed lubrication theory [venner_stokes_2003, marusic-paloka_second_2012, tavakol_extended_2017, sialmas_newtonian_2024] in the case of models based on perturbation theory. We assess the performance of these models at low Reynolds numbers by comparison to the Stokes equations. We propose a formulation of extended lubrication theory that builds on the model of Takeuchi & Gu (2019) [takeuchi_extended_2019]; our approach modifies the treatment of the velocity boundary conditions and ensures incompressibility. The pressure and velocity solutions from each model are compared in a variety of geometries: a logistic step, a backward facing step, and two versions of a triangular slider that we refer to as positive and negative based on the texture to be defined later on. Through these examples, we assess the sensitivity of each model to varying magnitudes of surface variation and discontinuities in the surface gradient in the low Reynolds number limit.
@@ 2.1 Lubrication Theory
We begin with a review of classical lubrication theory. This treatment may be found in various classical textbooks [leal_advanced_2007], and we include it here to establish notation and to make the presentation self-contained. The Navier-Stokes equations,
(1) $\frac{\partial p}{\partial x}$ $=\eta\Big(\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{2}u}{\partial y^{2}}\Big)-\rho\Big(\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}+v\frac{\partial u}{\partial y}\Big),$
(2) $\frac{\partial p}{\partial y}$ $=\eta\Big(\frac{\partial^{2}v}{\partial x^{2}}+\frac{\partial^{2}v}{\partial y^{2}}\Big)-\rho\Big(\frac{\partial v}{\partial t}+u\frac{\partial v}{\partial x}+v\frac{\partial v}{\partial y}\Big),$
@@
Lubrication theory provides a linearized approximation to the Navier-Stokes equations valid for geometries in which the fluid is constrained to a relatively thin domain. For the two-dimensional fluid domain $(x,y)\in[0,L_{x}]\times[0,h(x)]$, the characteristic length scales are $L_{x}>0$ and $L_{y}=\max h(x)>0$. Given a prescribed constant flux $\mathcal{Q}\neq 0$, the characteristic velocities are $U_{*}=\mathcal{Q}/L_{y}$ and $V_{*}=\mathcal{Q}/L_{x}$, the characteristic time is $T_{*}=L_{x}/U_{*}$, and the characteristic pressure is $P_{*}=\eta U_{*}L_{x}/L_{y}^{2}$. The dimensionless variables are $X=x/L_{x}$, $Y=y/L_{y}$, $H=h/L_{y}$, $U=u/U_{*}$, $V=v/V_{*}$, $T=t/T_{*}$ and $P=p/P_{*}$. The Reynolds number is $\text{Re}=\rho U_{*}L_{x}/\eta$ and the length scale ratio is $\varepsilon=L_{y}/L_{x}$. The Navier-Stokes equations Eqs. 1 and 2 in dimensionless terms are,
(4) $\frac{\partial P}{\partial X}$ $=\Big(\varepsilon^{2}\frac{\partial^{2}U}{\partial X^{2}}+\frac{\partial^{2}U}{\partial Y^{2}}\Big)-\varepsilon^{2}\text{Re}\Big(\frac{\partial U}{\partial T}+U\frac{\partial U}{\partial X}+V\frac{\partial U}{\partial Y}\Big),$
(5) $\frac{\partial P}{\partial Y}$ $=\varepsilon^{2}\Big(\varepsilon^{2}\frac{\partial^{2}V}{\partial X^{2}}+\frac{\partial^{2}V}{\partial Y^{2}}\Big)-\varepsilon^{4}\text{Re}\Big(\frac{\partial V}{\partial T}+U\frac{\partial V}{\partial X}+V\frac{\partial V}{\partial Y}\Big),$
@@
We also prescribe the fixed outlet pressure $P(1,0)=0$. Because the flux is prescribed, the pressure drop $\Delta P$ is given from the solution. Conversely, one may prescribe the pressure drop, in which case the flux would be determined by the solution. We define the dimensionless pressure drop,
(9) $\Delta P=\int_{0}^{H(0)}P(0,Y)\,dY-\int_{0}^{H(1)}P(1,Y)\,dY,$
and the dimensional pressure drop,
(10) $\Delta p=\frac{1}{L_{y}}\int_{0}^{h(0)}p(0,y)\,dy-\frac{1}{L_{y}}\int_{0}^{h(L_{x})}p(L_{x},y)\,dy,$
such that $\Delta P=\Delta p/P_{*}$.
@@
where $P_{0}$, $U_{0}$, $V_{0}$ denote the order $\varepsilon^{0}$ lubrication approximation. Together with incompressibility Eq. 3, Eqs. 11 and 12 constitute the governing equations for lubrication theory.
@@
The inlet pressure gradient Eq. 16 is obtained through integration of $U_{0}$ from Eq. 13 according to the flux condition Eq. 8, evaluated at $X=0$. Note that the flux appears in the scaling of the boundary velocity, $U_{b}=\mathcal{U}L_{y}/\mathcal{Q}$. Since the flux is prescribed, the inlet pressure $P_{0}(0)$ and the pressure drop $\Delta P_{0}$ are determined by the solution. The solution $P_{0}$ to the Reynolds equation is determined using a second-order accurate finite difference scheme, see Appendix A.
@@ 2.2 Perturbed Lubrication Theory
On substituting the ansatz Eqs. 17, 18 and 19 into Eqs. 4 and 5 at order $\varepsilon^{2}$, the velocity and pressure satisfy the following reduced equations,
@@
The order $\varepsilon^{0}$ Reynolds solution $U_{0}$, $V_{0}$, and $P_{0}$ satisfies the boundary conditions Eqs. 6, 7 and 16 respectively. The higher-order velocities $U$ and $V$ also satisfy Eqs. 6 and 7. The higher-order pressures $P$ satisfy $P(1,0)=0$, but do not necessarily satisfy the pressure gradient condition in Eq. 16. Instead, we require that the flux $\mathcal{Q}$ be the same as the flux of the order $\varepsilon^{0}$ solution, such that, in dimensionless terms, $\int_{0}^{H}U_{0}\,dY=1$ and $\int_{0}^{H}U_{k>0}\,dY=0$. Because the flux is prescribed, the average pressure drop $\Delta P_{\varepsilon^{k}\text{-PLT}}$ is determined by the solution. In general, the average pressure drops $\Delta P_{\varepsilon^{2}\text{-PLT}}=\Delta P_{0}+\varepsilon^{2}\Delta P_{2}$ and $\Delta P_{\varepsilon^{4}\text{-PLT}}=\Delta P_{0}+\varepsilon^{2}\Delta P_{2}+\varepsilon^{4}\Delta P_{4}$ will deviate from $\Delta P_{0}$. We determine $\Delta P_{k>0}$ through numerical integration according to Eq. 9.
@@
The method of perturbed solutions is also well suited to techniques such as diagonal Padé approximants or the Shanks’ transform [andrianov_pade_2021, housiadas_improved_2023, sialmas_newtonian_2024, sialmas_general_2025]. These techniques may be used to accelerate and extend the domain of the convergence for the series solution. Although such numerical methods are out of the scope of the present work, we consider this a potential avenue for future investigation and improvement of the PLT method.
@@ 3 Velocity-Adjusted Extended Lubrication Theory
Observe that if we assume $\varsigma=0$, the velocity field from Eqs. 28 and 29 is not necessarily incompressible, and Eq. 29 may not satisfy $V(X,H)=0$.
@@ 4.1 The logistic step
First, we consider the logistic step, shown in Fig. 1. The height function is given in dimensional terms by,
@@
The relative percent errors in $l_{2}$ norm of velocity and pressure compared with the Stokes solution are shown in Fig. 3. For the wide range of $\lambda<32$, the VA-ELT solution of pressure is significantly more accurate than the Reynolds and T.G.-ELT solutions, and the T.G.-ELT solution of pressure yields a marginal improvement over the Reynolds pressure. For $2<\lambda<8$, the $\varepsilon^{2}$-PLT solution is best at approximating the pressure, and the VA-ELT solution is best at approximating the velocity. The $\varepsilon^{2}$-PLT solution performs better than the $\varepsilon^{4}$-PLT solution for all $\lambda>2$. For $\lambda\geq 32$, the Reynolds solution has the smallest error in velocity and pressure; the Reynolds solution has the most stable error across the range of $\lambda$. Overall, both the $\varepsilon^{2}$-PLT and VA-ELT solutions perform well for the logistic step up to moderate surface variation. We note however, that all the lubrication models discussed here break down in the limit $\lambda\to\infty$ of a discontinuous step, see Appendix D.
@@ 4.2 The triangular slider
Next, we consider a triangular slider, shown in Fig. 4. The height function is given in dimensional terms by,
@@
The relative percent errors in the $l_{2}$ norm of pressure and velocity compared with the Stokes solution are shown in Fig. 7. The error in velocity is very similar for all of the models. For pressure, the VA-ELT solution is a consistent improvement compared to the Reynolds and T.G.-ELT solutions, and the $\varepsilon^{k}$-PLT solutions perform better still. For negative texturing ($H_{v}<1$), the $\varepsilon^{k}$-PLT solutions of pressure have significantly smaller error relative to both the VA-ELT and Reynolds solutions; for positive texturing ($H_{v}>1$), the $\varepsilon^{k}$-PLT and VA-ELT solutions are more similar. The error in both pressure and velocity decreases as the difference between the minimum and maximum height decreases. In particular, for the case of negative texturing, the increasing magnitude of surface gradient as the minimum height decreases corresponds to an increase in error for pressure and velocity; this emphasizes the importance of considering the magnitude of surface gradients as well as the length scale ratio when selecting a fluid model.
@@ Appendix A The Reynolds equation finite difference solution
Define the uniform discretisation of $[0,1]$,
@@
Appendix B The Stokes Equation
The Stokes equation is derived from the Navier-Stokes equation under the assumption of zero Reynolds number and a steady state flow [leal_advanced_2007]. Compared with the lubrication approximation, Stokes flow is formulated with a single characteristic velocity $U_{*}$ and length scale $L_{x}$. The characteristic pressure is now $\tilde{P}_{*}=\eta U_{*}/L_{x}$. Consequently, Stokes flow does not involve the length scale ratio $\varepsilon$ and is not sensitive to surface variation as the lubrication models are.
With the dimensionless variables $X=x/L_{x}$, $Y=y/L_{x}$, $U=u/U_{*}$, $V=v/U_{*}$, $T=t/T_{*}$ and $P=p/\tilde{P}_{*}$, the Navier-Stokes equations Eqs. 1 and 2 may be expressed as
(35) $\frac{\partial P}{\partial X}$ $=\Big(\frac{\partial^{2}U}{\partial X^{2}}+\frac{\partial^{2}U}{\partial Y^{2}}\Big)-\text{Re}\Big(\frac{\partial U}{\partial T}+U\frac{\partial U}{\partial X}+V\frac{\partial U}{\partial Y}\Big),$
(36) $\frac{\partial P}{\partial Y}$ $=\Big(\frac{\partial^{2}V}{\partial X^{2}}+\frac{\partial^{2}V}{\partial Y^{2}}\Big)-\text{Re}\Big(\frac{\partial V}{\partial T}+U\frac{\partial V}{\partial X}+V\frac{\partial V}{\partial Y}\Big).$